OpenStudy (anonymous):

(fifth root(5x^2y)^3

jhonyy9 (jhonyy9):

\[\sqrt[5](5x^2y)^3 =\]

jhonyy9 (jhonyy9):

you have thought it so ?

OpenStudy (anonymous):

yes. root sign over all of (5x^2y)

jhonyy9 (jhonyy9):

but there is (5x^2y)^3 cubed the paraentheses yes ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

so can you calcule it ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

can you show me how?

OpenStudy (anonymous):

Please?

jhonyy9 (jhonyy9):

so check it will be very easy will be fifthroot5^3 *fifthrootx^6 *fifthrooty^3 = can you continue it ?

OpenStudy (anonymous):

?

jhonyy9 (jhonyy9):

so from this just fifthrootx^6 can simplifie again like fifthroot(x^5*x) = xfifthrootx ok ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

so can you rewrite it for the right finaly answer ?

OpenStudy (anonymous):

no

jhonyy9 (jhonyy9):

so will be x*fifthroot(x(5y)^3)

jhonyy9 (jhonyy9):

so do you you understand it now ?

OpenStudy (anonymous):

Yes. but none of the multiple choice answers on my homework fit

OpenStudy (anonymous):

(5xy)^6/5, (5x^2y)^5/3 (5xy5/6) (5x^2y)^15 (5x^2y)^3/5

jhonyy9 (jhonyy9):

so the last one is the right answer because if you revrite an equation inside radical so than the denominator of the fraction from exponent mean the indice of radical ok ?

OpenStudy (anonymous):

I thought that (5x^2y)3/5 was right but then It didn't fit my calculations.

jhonyy9 (jhonyy9):

but i think that the last one is right sure

OpenStudy (anonymous):

The one that I thought was right?

jhonyy9 (jhonyy9):

so than yes if you have thought it in this way

OpenStudy (anonymous):

ok. thank you. I am never going to make it through math and my math teacher doesn't explain anything

jhonyy9 (jhonyy9):

so was my pleasure to help you good luck bye

OpenStudy (anonymous):

Bye